I am trying to figure out this problem!!! It is
"The length of a rectangle is 4 units less than 3 times the width. The perimeter is 22 units more than twice the width. Find the length and width of the rectangle."
i have already tried X=4+(3w)+(w+22) but I don't know if it is right, or has been formatted correctly.
Thank You for your Time,
Mr_Bunston

Answer by ewatrrr(24785)

 
Hi,
.
2(3w-4) + 2w = 2w+22
6w - 8 + 2w = 2w + 22
6w = 30
w = 5 units and L = 11 units
checking...
32 = 10 + 22